Walt made an extra $5000 last year from a part time job. He invested part of the money at 8% and the rest at 9%. He made a total of $440 in interest. How much was invested at 9%?
Answer by mananth(16949)

Part I 8.00% per annum ------------- Amount invested =x
Part II 9.00% per annum ------------ Amount invested = y
5000
Interest----- 440

Part I 8.00% per annum ---x
Part II 9.00% per annum ---y
Total investment
x + 1 y= 5000 -------------1
Interest on both investments
8.00% x + 9.00% y= 440
Multiply by 2
8 x + 9 y= 44000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -40000.00
Add this to (2)
0 x 1 y= 4000
divide by 1
y = 4000
Part I 8.00% $ 1000
Part II 9.00% $ 4000

CHECK
1000 --------- 8.00% ------- 80.00
4000 ------------- 9.00% ------- 360.00
Total -------------------- 440.00
$4000 @ 9%

m.ananth@hotmail.ca